If you deposit money today in an account that pays 5.7% annual interest, how long will it take to double your money? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value($2x)
P=present value($x say)
r=rate of interest
n=time period.

2x=x*(1.057)^n

2=(1.057)^n

Taking log on both sides;

log 2=n*log (1.057)

n=log 2/log (1.057)

=12.50 years(Approx).